OpenStudy (anonymous):

evaluate sec(sin^-1 4/5)

OpenStudy (anonymous):

Suppose \(\theta=\sin^{-1}\dfrac{4}{5}\), so that \(\sin\theta=\dfrac{4}{5}\). This gives you a triangle:  What's the length of the missing side? If \(\theta\) is the angle on the left, what would its secant be?

OpenStudy (anonymous):

sin of -4/5

OpenStudy (anonymous):

Um, what?

OpenStudy (anonymous):

Guessing I am not really that good with this trig stuff

OpenStudy (anonymous):

 From the Pythagorean theorem, you get \[\begin{align*}x^2+4^2&=5^2\\ x^2&=25-16\\ x^2&=9\\ x&=3 \end{align*}\] So if \(\cos\theta=\dfrac{\text{adjacent}}{\text{hypotenuse}}\) and \(\sec\theta=\dfrac{1}{\cos\theta}\), what do you get?

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

Correct